I need help with this homework, and you guys have never failed me!
Graph each linear equation by finding and plotting its intercepts.
My teacher wants three different plots, I used (0,3) and (3,0) can't figure out the next.

Answer by shree840(260)

I do not understand how you got (3,0) and (0,3)
I think your teacher wants 3 diff examples
let us say 1. eq is 3x+y=5 y intercept when x=0 y=5 (0,5)
x intercept when y=0 x+5/3 (5/3,0)
eq 2 2y-3x=6 y intercept when x=0 y=3 (0,3)
x Intercept when y=0 x=-2 (-2,0)
similarly take eq 3